find the equation of the line that contains (-8,-1)and is parallel to the line x-10y=5
Answer by mananth(16946)

1 x + -10 y = 5
Find the slope of this line
make y the subject
-10 y =-1x+5
Divide by -10
y = 0.1x-0.5
Compare this equation with y=mx+b
slope m = 0.1
The slope of a line parallel to the above line will be the same
The slope of the required line will be 0.1
m= 0.1 ,point ( -8 , -1 )
Find b by plugging the values of m & the point in
y=mx+b
-1 = -0.80 + b
b= -0.2
m= 0.1
Plug value of the slope and b in y = mx +b
The required equation is y = 0.1x-0.2
(y = 1/10 x - 1/ 5)
m.ananth@hotmail.ca